A class has a pinic and eats for 1/2 an hour,1 and 3/4 of an hour for playing,1 and 2/3 of more playing of an hour, eats desert for 1/5 of an hour, and spends 1/20 of an hour running to get out of the rain. if the picnic started at 11 am when did it end ? How much more time was spent playing than eating ? Please answer these quesions for me and explain it well step by step. thankyou.
1
Expert's answer
2012-12-07T03:42:10-0500
They started at 11.00 and ate for eats for 1/2 an hour( 1/2 hour =30 min)then playing(1 and 3/4 =1.45 hour). At that moment time was
11.00+00.30+1.00+00.45 = 13.15
Thenthey again played about (1 and 2/3=1.40) (14.55 pm) then they spent( 1/5
of an hour = 12 min) 12 min eating desert
Atthat moment time is 13.15+1.00+0.40+0.12 =15.07.Then they spent(1/20 of an
hour = 3 min) 3 min to running to get out of the rain.
So at 15.10 pm they ended.
Totally they spent 1.45+1.40 = 2 hour 25 min for playing